CH01: Homework Questions 3696 Fullscreen Assigned as Homework O Homework Question 1.12 Fill in the Blanks Predict the atomic mass and write the shorthand that describes these isotopes. For example, hydrogen with 2 neutrons has atomic mass 3 and shorthand H-3 Beryllium with 4 neutrons: atbmic mass ; shorthand Sodium with 17 neutrons: atomic mass ; shorthand Silicon with 10 neutrons: atomic mass shorthand: Chlorine with 14 neutrons: atomic mass ; shorthand Cobalt with 33 neutrons: atomic mass ; shorthand Strontium with 52 neutrons: atomic mass ; shorthand: r 1 Fullsc Unanswered 3 attempts left
Answer #1

Atomic Mass = No.of neutrons + No.of protons

where, No.of protons = Atomic number

a)

Beryllium atomic number = No.of protons in Beryllium atom = 4

and given, no.of neutrons = 4

Therefore,

Atomic mass = no.of neutrons + No.of protons = 4+4 = 8 and short hand is Be-8

Therefore, Answer is: Atomic mass = 8 and shorthand = Be-8

Similarly,

b)

Given, No.of neutrons in Na atom = 17

Atomic number of Na = No.of protons = 11

Therefore, Atomic mass = 17+11 =28 and short hand = Sodium-28

Answer: Atomic mass of Sodium =28 and short hand = Sodium-28

c)

No.of neutrons of in the given Silicon atom = 10 neutrons

Atomic number of Silicon = 14

Therefore, Atomic mass = 10+14 =24 and short hand = Silicon-24

d)

No.of neutrons of in the given Chlorine atom = 14 neutrons

Atomic number of Chlorine = 17

Therefore, Atomic mass = 14+17 =31 and short hand = Chlorine-31

e)

No.of neutrons of in the given Cobalt atom = 33 neutrons

Atomic number of Cobalt =27

Therefore, Atomic mass = 33+27 =60 and short hand = Cobalt-60

f)

No.of neutrons of in the given Strontium atom = 52 neutrons

Atomic number of Strontium = 38

Therefore, Atomic mass = 52+38 = 90 and short hand = Strontium-90